# Boosting drug-disease association prediction for drug repositioning via dual-feature extraction and cross-dual-domain decoding

### v3

@@ Abstract
Index Terms:
Drug Repositioning, Self-Attention Mechanism, Dual-Feature Extraction, Cross-Dual-Domain
I Introduction
@@
Drug development is a significant yet complex and time-consuming process. It consists of three main stages: the discovery stage, the preclinical stage, and the clinical stage. Typically, the entire process takes about 10 to 15 years and costs around \$2.8 billion on average [1, 2]. Despite ongoing efforts and innovations, the traditional drug development model faces considerable risks, often resulting in project terminations due to concerns about drug efficacy, safety, or commercial viability. Notably, a high percentage (80-90%) of these projects fail during clinical trials. Consequently, finding a more efficient and precise method for drug development is critically important and valuable.
@@
Given the high cost of drug development, computational drug repositioning has become a powerful strategy for discovering alternative uses of already approved drugs while considering their potential effects [3]. This approach utilizes computational and bioinformatics tools, providing greater efficiency and cost-effectiveness compared to traditional drug research and development methods. It enables the rapid screening of numerous candidate drugs and accelerates the development of new therapeutic options. Repurposing existing drugs can significantly lower development costs when compared to creating new drugs from scratch. In response to health crises, such as pandemics, repurposing known drugs offers a quicker solution than developing entirely new treatments. Furthermore, this method can unveil new applications for drugs that may be effective against diseases that currently lack effective therapies, thereby addressing major health challenges.
@@
Machine learning has been commonly used in computational drug repositioning by information extraction from biomedical data. Techniques like random forest (RF), support vector machine (SVM), and decision trees are often applied for this purpose [4]. For instance, [5] introduced a novel drug-target interaction screening framework called PUDTI, which utilized SVM to identify potential drug repositioning targets. [6] proposed a method for predicting drug-target interactions using Lasso dimensionality reduction and a random forest classifier. [7] employed random forests in their final prediction step, while other techniques such as decision trees [8], matrix factorization[9, 10], and logistic regression [11] have also been developed.
@@
In recent years, deep learning technology has become a powerful tool in bioinformatics, as evidenced by [12, 13, 14, 15]. Unlike traditional machine learning methods, deep learning techniques leverage deep architectures to effectively capture high-level latent representations without requiring manual feature selection and tuning [16]. Existing deep learning methods for drug repositioning often incorporate graph neural networks like Graph Convolutional Networks (GCN) and Graph Attention Networks (GAT) to for model building [17, 18, 19, 20, 21]. For instance, [17] performed graph convolution operations on the drug-disease heterogeneous network to encode drugs and diseases but did not account for the differences in encoding from different network spaces. [18] distinguished intra-domain and inter-domain feature extraction on the drug-disease heterogeneous network for encoding drugs and diseases. [21] extracted features of drugs and diseases from multiple neighborhood spaces and designed a contrastive learning loss to fuse these features. However, these studies did not consider calculating information weights from different neighbors. Authors in [22, 23] used graph attention networks (GAT) to calculate the weights of neighbor aggregation and extract key neighbor information for encoding. Most deep learning methods have primarily focused on extracting valuable features from neighborhoods. However, these approaches often rely on a single encoding method, which does not intentionally account for relation prediction. As a result, a significant amount of noise can arise during model training, hindering the effective extraction of drug-disease association information and ultimately impacting the accuracy of the prediction results.
@@
The rest of this paper is structured as follows: Section II introduces the datasets and methods, providing details on the DFDRNN implementation; Section III explains the experimental setup and discusses the results; Section IV concludes the study and suggests future research directions.
II Materials and Methods
II-A Datasets
We performed experiments on four benchmark datasets. The first, Gdataset [24], comprises 593 drugs from DrugBank [25] and 313 diseases from the Online Mendelian Inheritance in Man (OMIM) [26], with a total of 1933 drug-disease associations. The second dataset, Cdataset [27], consists of 663 drugs listed in DrugBank, 409 diseases from OMIM, and 2532 known associations. Ldataset [17], the third dataset, contains 18,416 associations between 269 drugs and 598 diseases, derived from the Comparative Toxicogenomics Database (CTD) [28]. Lastly, LRSSL [29] contains 763 drugs, 681 diseases, and 3051 validated drug-disease associations. A summary of the statistics for these datasets can be found in Table I.
This study determines drug similarity based on chemical structures obtained from DrugBank using standardized SMILES notation [30]. For disease similarity, it is determined based on disease phenotypes and computed via MimMiner [31], which measures similarity by analyzing the MeSH terms in the medical descriptions of diseases found in the OMIM database [32].
TABLE I: Summary statistics of the four datasets
@@ II-B1 Three key networks
II-B2 Self-attention mechanism
The self-attention mechanism (SAM) can be described as a process associated with three vectors, i.e., one query vector $Q_{i}$, some key vectors $K_{j}$, and some value vector $V_{j}$, which maps $Q_{i}$ and a group of $K_{j}$-$V_{j}$ pairs to a new output vector. The output of the SAM is a weighted sum of $V_{j}$ vectors, where the weight for each $V_{j}$ is computed by taking the dot product of the $Q_{i}$ with the corresponding $K_{j}$ [33]. In our SDDFE and CDDFE modules, we use the multi-head SAM to aggregate neighbor information for drugs and diseases. We use the notation $\mathbf{sam}$ to represent the multi-head SAM mapping that involves two parameters $E$ and $H^{in}$, where $E\in\{E_{s},E_{a}\}$ and $H^{in}\in\mathbb{R}^{(n+m)\times k}$ is an input feature matrix with embedding dimension $k$. The specific implementation of $\mathbf{sam}$ can be described as follows:
@@ II-B2 Self-attention mechanism
After the SAM function, a fully connected layer is employed to further extract the features of drugs and diseases. We use the notation $\mathbf{samf}$ to denote the whole process, i.e., $\mathbf{samf}$ is the combination of $\mathbf{sam}$ with a fully connected (FC) layer. Let $w_{r}\in\mathbb{R}^{k\times k}$ and $w_{d}\in\mathbb{R}^{k\times k}$ be trainable parameter matrices, $b_{r}$ and $b_{d}$ be biases, and $\sigma$ be a LeakyRELU [34] activation function. Then,
@@ II-B4 Decoder
6:        for $\ell=$ to $L-1$
7:              Compute SDDFE $\hat{H}^{\ell+1}_{s},\hat{H}^{\ell+1}_{a}$ with Equation (10).
8:              Compute CDDFE $\tilde{H}^{\ell+1}_{s},\tilde{H}^{\ell+1}_{a}$ with Equation (11).
9:              Fuse the same feature and the input feature yields $H^{\ell+1}_{s}$ and $H^{\ell+1}_{a}$ with Equation (12) (13)
10:        end for
11:        Utilizing layer attention mechanism to compute the final embedding with Equation (14)
12:        Obtain Computer the prediction score matrix $s$ with Equation (15)(16)(17)
13:        Update the parameters $w^{Q},w^{K},w^{V},w_{r},w_{b},\ldots$ with Equation (18).
14:        $c\leftarrow c+1$.
15:   end for
II-C Optimization
@@
The DFDRNN model is optimized using the Adam optimizer [35], with parameters initialized through Xavier initialization [36]. To enhance generalization, we randomly drop edges and apply regular dropout techniques [37].
III Experiments and Results
III-A Baseline methods
@@
$\bullet$ LAGCN [17]: This method utilizes multi-layer GCNs for neighbor information propagation on heterogeneous networks and introduces the layer attention mechanism.
$\bullet$ DRHGCN [18]: This model uses GCN and bilinear aggregators to design inter-domain and intra-domain embeddings for predicting drug-disease associations.
$\bullet$ DRWBNCF [38]: This model focuses on the interaction information between neighbors and uses generalized matrix factorization collaborative filtering to predict drug-disease associations.
$\bullet$ NCH-DDA [21]: This method involves the design of a single-neighborhood feature extraction module and a multi-neighborhood feature extraction module to extract features for drugs and diseases. Additionally, it introduces neighborhood contrastive learning for drug-disease association prediction.
$\bullet$ HDGAT [23]: This approach combines graph convolutional neural networks with bidirectional extended short-term memory networks and integrates layer attention mechanisms and residual connections.
$\bullet$ DRGBCN [20]: This method utilizes layer attention graph convolutional networks to encode drugs and diseases and employs bilinear attention network modules to enhance their intricate relationship.
III-B Experimental parameter configuration
@@
TABLE II: AUROC and AUPR results for DFDRNN and comparative models using 10-times repeated 10-fold cross-validation
@@
III-C Evaluation metrics
We classify the prediction outcomes into four categories. As illustrated in Table III, TP denotes true positives that are correctly identified as positive, FN represents false negatives that are wrongly classified as negative, FP indicates false positives that are incorrectly classified as positive, and TN stands for true negatives that are accurately identified as negative. Based on these definitions, we proceed to calculate the True Positive Rate (TPR or Recall), False Positive Rate (FPR), and Precision using the following formulas:
TABLE III: The statistics of the four categories
@@ III-C Evaluation metrics
III-D DFDRNN’s performance in cross-validation
We utilize 10-fold cross-validation (10-CV) to evaluate the model’s performance. The dataset is randomly divided into ten distinct subsets, each comprising 10% known drug-disease associations and 10% unknown pairs. In each iteration, one subset is used as the test set, while the other nine serve as the training set. This process is repeated ten times with varying random seeds, allowing for the calculation of the average and standard deviation of the performance metrics. The experimental outcomes, including AUROC and AUPR scores, are summarized in Table II, while the corresponding ROC and PR curves are provided in Supplementary Figures S1-S4.
In Table II, it is evident that DFDRNN outperformed all other models by achieving the highest average AUROC of 0.946 and the highest average AUPR of 0.597 across the four datasets. This performance surpassed the second-ranked NCH-DDA model by 0.7% and 1.8% in AUROC and AUPR, respectively. DFDRNN demonstrated the best AUROC and AUPR on all four datasets, showcasing its predictive solid capability for potential drug-disease associations based on known associations. It’s worth mentioning that NCH-DDA performed exceptionally well, coming in second only to DFDRNN, which may be attributed to its feature extraction from different neighborhoods and the fusion of features using contrastive learning loss.
III-E Effect of the number of nearest neighbors
@@
III-F Ablation analysis
III-F1 Encoder
@@
Table IV presents the AUROC and AUPR values of DFDRNN and its four variants under 10-CV. By comparing the two variants of DFDRNN, DFDRNN-SF and DFDRNN-AF, we observe that DFDRNN-SF performs better than DFDRNN-AF on the Gdataset, Cdataset, and LRSSL, indicating the contribution of the similarity feature is greater than that of the association feature on these three datasets. However, on the Ldataset, the contribution of the association feature is greater than that of the similarity feature. Compared to DFDRNN, simultaneously encoding both features can significantly improve the model’s performance. Compared to DFDRNN, DFDRNN-GCN demonstrates a noticeable decrease in AUROC values, highlighting the strong performance of SAM in extracting neighbor information. Furthermore, compared to DFDRNN, DFDRNN-noCF shows a significant decrease in AUPR on the Gdataset, Cdataset, and LRSSL datasets, and performs the worst on the Ldataset, indicating the necessity of changing features in the CDDFE module.
TABLE IV: The performance of DFDRNN and its four variants
@@ III-F2 Decoder
III-G Identification of candidates for new diseases
In late 2019, the discovery of COVID-19 presented a challenge in promptly researching and developing suitable drugs for treatment. Repurposing already approved drugs, such as remdesivir, favipiravir, and ribavirin, played a significant role in the timely control of the epidemic [39]. Finding suitable drugs for emerging diseases is a complex and crucial task. To assess DFDRNN’s ability to predict candidate drugs for new diseases, we applied leave-one-out cross-validation (LOOCV) using the Gdataset. For each disease $d_{x}$, all drug-disease associations linked to $d_{x}$ were removed and treated as test data, while the remaining associations served as training data. The prediction relied on the similarity between $d_{x}$ and other diseases to identify its candidate drugs.
@@ III-G Identification of candidates for new diseases
III-H Case study
To evaluate the practical effectiveness of DFDRNN, we performed case studies on two diseases, Alzheimer’s disease (AD) and Parkinson’s disease (PD), from the Gdataset. Using all drug-disease associations within the Gdataset as the training data, we trained DFDRNN to predict association probabilities for all unknown drug-disease pairs. We then ranked the drugs in descending order of predicted probabilities, identifying the top 10 candidate drugs for AD and PD. To validate DFDRNN’s predictions, we referenced reliable sources and clinical trials, including DrugBank [25], PubChem [40], CTD [28], DrugCentral [41], and ClinicalTrials.
AD, the most prevalent form of progressive dementia in the elderly, is a neurodegenerative disorder. Table V displays the top 10 drugs predicted by DFDRNN for AD. Among these, Amantadine is the first drug predicted to hold therapeutic potential for AD. Amantadine, commonly used as an antiviral and antiparkinson drug, exhibits actions as an antiviral, antiparkinson, dopaminergic, and analgesic agent. It can also be utilized for the prevention or symptomatic treatment of influenza A. Previous studies have shown improvements in the mental state of two AD patients following Amantadine treatment [42]. Carbamazepine functions as an anticonvulsant and analgesic drug. Earlier studies have suggested that the combined therapy of carbamazepine and haloperidol holds promise in the clinical management of elderly AD patients [43]. DFDRNN predicts an association with carbamazepine, which has been validated by CTD. Moreover, all the other drugs listed in Table V have been confirmed by relevant authoritative data sources, with a success rate of 100%.
TABLE V: The top 10 candidate drugs for AD are ranked based on the DFDRNN prediction score

### v4

@@ Abstract
keywords:
American Chemical Society,
[…]
IR,NMR,UV
1 Introduction
@@
Drug development is a significant yet complex and time-consuming process. It consists of three main stages: the discovery stage, the preclinical stage, and the clinical stage. Typically, the entire process takes about 10 to 15 years and costs around \$2.8 billion on average 1, 2. Despite ongoing efforts and innovations, the traditional drug development model faces considerable risks, often resulting in project terminations due to concerns about drug efficacy, safety, or commercial viability. Notably, a high percentage (80-90%) of these projects fail during clinical trials. Consequently, finding a more efficient and precise method for drug development is critically important and valuable.
@@
Given the high cost of drug development, computational drug repositioning has become a powerful strategy for discovering alternative uses of already approved drugs while considering their potential effects 3. This approach utilizes computational and bioinformatics tools, providing greater efficiency and cost-effectiveness compared to traditional drug research and development methods. It enables the rapid screening of numerous candidate drugs and accelerates the development of new therapeutic options. Repurposing existing drugs can significantly lower development costs when compared to creating new drugs from scratch. In response to health crises, such as pandemics, repurposing known drugs offers a quicker solution than developing entirely new treatments. Furthermore, this method can unveil new applications for drugs that may be effective against diseases that currently lack effective therapies, thereby addressing major health challenges.
@@
Machine learning has been commonly used in computational drug repositioning by information extraction from biomedical data. Techniques like random forest (RF), support vector machine (SVM), and decision trees are often applied for this purpose 4. For instance, 5 introduced a novel drug-target interaction screening framework called PUDTI, which utilized SVM to identify potential drug repositioning targets. 6 proposed a method for predicting drug-target interactions using Lasso dimensionality reduction and a random forest classifier. 7 employed random forests in their final prediction step, while other techniques such as decision trees 8, matrix factorization9, 10, 11, and logistic regression 12 have also been developed.
@@
In recent years, deep learning technology has become a powerful tool in bioinformatics, as evidenced by 13, 14, 15, 16. Unlike traditional machine learning methods, deep learning techniques leverage deep architectures to effectively capture high-level latent representations without requiring manual feature selection and tuning 17. Existing deep learning methods for drug repositioning often incorporate graph neural networks like Graph Convolutional Networks (GCN) and Graph Attention Networks (GAT) to for model building 18, 19, 20, 21, 22. For instance, 18 performed graph convolution operations on the drug-disease heterogeneous network to encode drugs and diseases but did not account for the differences in encoding from different network spaces. 19 distinguished intra-domain and inter-domain feature extraction on the drug-disease heterogeneous network for encoding drugs and diseases. 22 extracted features of drugs and diseases from multiple neighborhood spaces and designed a contrastive learning loss to fuse these features. However, these studies did not consider calculating information weights from different neighbors. Authors in 23, 24, 25 used graph attention networks (GAT) to calculate the weights of neighbor aggregation and extract key neighbor information for encoding. Most deep learning methods have primarily focused on extracting valuable features from neighborhoods. However, these approaches often rely on a single encoding method, which does not intentionally account for relation prediction. As a result, a significant amount of noise can arise during model training, hindering the effective extraction of drug-disease association information and ultimately impacting the accuracy of the prediction results.
@@
The rest of this paper is structured as follows: Section 2 introduces the datasets and methods, providing details on the DFDRNN implementation; Section 3 explains the experimental setup and discusses the results; Section 4 concludes the study and suggests future research directions.
2 Materials and Methods
2.1 Datasets
We performed experiments on four benchmark datasets. The first, Gdataset 26, comprises 593 drugs from DrugBank 27 and 313 diseases from the Online Mendelian Inheritance in Man (OMIM) 28, with a total of 1933 drug-disease associations. The second dataset, Cdataset 29, consists of 663 drugs listed in DrugBank, 409 diseases from OMIM, and 2532 known associations. Ldataset 18, the third dataset, contains 18,416 associations between 269 drugs and 598 diseases, derived from the Comparative Toxicogenomics Database (CTD) 30. Lastly, LRSSL 31 contains 763 drugs, 681 diseases, and 3051 validated drug-disease associations. A summary of the statistics for these datasets can be found in Table 1.
This study determines drug similarity based on chemical structures obtained from DrugBank using standardized SMILES notation 32. For disease similarity, it is determined based on disease phenotypes and computed via MimMiner 33, which measures similarity by analyzing the MeSH terms in the medical descriptions of diseases found in the OMIM database 34.
Table 1: Summary statistics of the four datasets
@@ 2.2.1 Three key networks
2.2.2 Self-attention mechanism
The self-attention mechanism (SAM) can be described as a process associated with three vectors, i.e., one query vector $Q_{i}$, some key vectors $K_{j}$, and some value vector $V_{j}$, which maps $Q_{i}$ and a group of $K_{j}$-$V_{j}$ pairs to a new output vector. The output of the SAM is a weighted sum of $V_{j}$ vectors, where the weight for each $V_{j}$ is computed by taking the dot product of the $Q_{i}$ with the corresponding $K_{j}$ 35. In our SDDFE and CDDFE modules, we use the multi-head SAM to aggregate neighbor information for drugs and diseases. We use the notation $\mathbf{sam}$ to represent the multi-head SAM mapping that involves two parameters $E$ and $H^{in}$, where $E\in\{E_{s},E_{a}\}$ and $H^{in}\in\mathbb{R}^{(n+m)\times k}$ is an input feature matrix with embedding dimension $k$. The specific implementation of $\mathbf{sam}$ can be described as follows:
@@ 2.2.2 Self-attention mechanism
After the SAM function, a fully connected layer is employed to further extract the features of drugs and diseases. We use the notation $\mathbf{samf}$ to denote the whole process, i.e., $\mathbf{samf}$ is the combination of $\mathbf{sam}$ with a fully connected (FC) layer. Let $w_{r}\in\mathbb{R}^{k\times k}$ and $w_{d}\in\mathbb{R}^{k\times k}$ be trainable parameter matrices, $b_{r}$ and $b_{d}$ be biases, and $\sigma$ be a LeakyRELU 36 activation function. Then,
@@ 2.2.4 Decoder
6:
7:     for $\ell=$ to $L-1$ do
8:                    Compute SDDFE $\hat{H}^{\ell+1}_{s},\hat{H}^{\ell+1}_{a}$ with Equation (10).
9:                    Compute CDDFE $\tilde{H}^{\ell+1}_{s},\tilde{H}^{\ell+1}_{a}$ with Equation (11).
10:                    Fusing the same feature with Equations (12) and (13) produces $H^{\ell+1}_{s}$ and $H^{\ell+1}_{a}$.
11:
12:     end for
13:           Final embedding via layer attention and Equation (14).
14:           Obtain Computer the prediction score matrix $s$ with Equation (15)(16)(17)
15:           Update the parameters $w^{Q},w^{K},w^{V},w_{r},w_{b},\ldots$ with Equation (18).
16:           $c\leftarrow c+1$.
17:  end for
2.3 Optimization
@@
The DFDRNN model is optimized using the Adam optimizer 37, with parameters initialized through Xavier initialization 38. To enhance generalization, we randomly drop edges and apply regular dropout techniques 39.
3 Experiments and Results
3.1 Baseline methods
@@
$\bullet$ LAGCN 18: This method utilizes multi-layer GCNs for neighbor information propagation on heterogeneous networks and introduces the layer attention mechanism.
$\bullet$ DRHGCN 19: This model uses GCN and bilinear aggregators to design inter-domain and intra-domain embeddings for predicting drug-disease associations.
$\bullet$ DRWBNCF 40: This model focuses on the interaction information between neighbors and uses generalized matrix factorization collaborative filtering to predict drug-disease associations.
$\bullet$ NCH-DDA 22: This method involves the design of a single-neighborhood feature extraction module and a multi-neighborhood feature extraction module to extract features for drugs and diseases. Additionally, it introduces neighborhood contrastive learning for drug-disease association prediction.
$\bullet$ HDGAT 24: This approach combines graph convolutional neural networks with bidirectional extended short-term memory networks and integrates layer attention mechanisms and residual connections.
$\bullet$ DRGBCN 21: This method utilizes layer attention graph convolutional networks to encode drugs and diseases and employs bilinear attention network modules to enhance their intricate relationship.
3.2 Experimental parameter configuration
@@
Table 2: AUROC and AUPR results for DFDRNN and comparative models using 10-times repeated 10-fold cross-validation
@@
3.3 Evaluation metrics
We classify the prediction outcomes into four categories. As illustrated in Table 3, TP denotes true positives that are correctly identified as positive, FN represents false negatives that are wrongly classified as negative, FP indicates false positives that are incorrectly classified as positive, and TN stands for true negatives that are accurately identified as negative. Based on these definitions, we proceed to calculate the True Positive Rate (TPR or Recall), False Positive Rate (FPR), and Precision using the following formulas:
Table 3: The statistics of the four categories
@@ 3.3 Evaluation metrics
3.4 DFDRNN’s performance in cross-validation
We utilize 10-fold cross-validation (10-CV) to evaluate the model’s performance. The dataset is randomly divided into ten distinct subsets, each comprising 10% known drug-disease associations and 10% unknown pairs. In each iteration, one subset is used as the test set, while the other nine serve as the training set. This process is repeated ten times with varying random seeds, allowing for the calculation of the average and standard deviation of the performance metrics. The experimental outcomes, including AUROC and AUPR scores, are summarized in Table 2, while the corresponding ROC and PR curves are provided in Supplementary Figures S1-S4.
In Table 2, it is evident that DFDRNN outperformed all other models by achieving the highest average AUROC of 0.946 and the highest average AUPR of 0.597 across the four datasets. This performance surpassed the second-ranked NCH-DDA model by 0.7% and 1.8% in AUROC and AUPR, respectively. DFDRNN demonstrated the best AUROC and AUPR on all four datasets, showcasing its predictive solid capability for potential drug-disease associations based on known associations. It’s worth mentioning that NCH-DDA performed exceptionally well, coming in second only to DFDRNN, which may be attributed to its feature extraction from different neighborhoods and the fusion of features using contrastive learning loss.
3.5 Effect of the number of nearest neighbors
@@
3.6 Ablation analysis
3.6.1 Encoder
@@
Table 4 presents the AUROC and AUPR values of DFDRNN and its four variants under 10-CV. By comparing the two variants of DFDRNN, DFDRNN-SF and DFDRNN-AF, we observe that DFDRNN-SF performs better than DFDRNN-AF on the Gdataset, Cdataset, and LRSSL, indicating the contribution of the similarity feature is greater than that of the association feature on these three datasets. However, on the Ldataset, the contribution of the association feature is greater than that of the similarity feature. Compared to DFDRNN, simultaneously encoding both features can significantly improve the model’s performance. Compared to DFDRNN, DFDRNN-GCN demonstrates a noticeable decrease in AUROC values, highlighting the strong performance of SAM in extracting neighbor information. Furthermore, compared to DFDRNN, DFDRNN-noCF shows a significant decrease in AUPR on the Gdataset, Cdataset, and LRSSL datasets, and performs the worst on the Ldataset, indicating the necessity of changing features in the CDDFE module.
Table 4: The performance of DFDRNN and its four variants
@@ 3.6.2 Decoder
3.7 Identification of candidates for new diseases
In late 2019, the discovery of COVID-19 presented a challenge in promptly researching and developing suitable drugs for treatment. Repurposing already approved drugs, such as remdesivir, favipiravir, and ribavirin, played a significant role in the timely control of the epidemic 41. Finding suitable drugs for emerging diseases is a complex and crucial task. To assess DFDRNN’s ability to predict candidate drugs for new diseases, we applied leave-one-out cross-validation (LOOCV) using the Gdataset. For each disease $d_{x}$, all drug-disease associations linked to $d_{x}$ were removed and treated as test data, while the remaining associations served as training data. The prediction relied on the similarity between $d_{x}$ and other diseases to identify its candidate drugs.
@@ 3.7 Identification of candidates for new diseases
3.8 Case study
To evaluate the practical effectiveness of DFDRNN, we performed case studies on two diseases, Alzheimer’s disease (AD) and Parkinson’s disease (PD), from the Gdataset. Using all drug-disease associations within the Gdataset as the training data, we trained DFDRNN to predict association probabilities for all unknown drug-disease pairs. We then ranked the drugs in descending order of predicted probabilities, identifying the top 10 candidate drugs for AD and PD. To validate DFDRNN’s predictions, we referenced reliable sources and clinical trials, including DrugBank 27, PubChem 42, CTD 30, DrugCentral 43, and ClinicalTrials.
AD, the most prevalent form of progressive dementia in the elderly, is a neurodegenerative disorder. Table 5 displays the top 10 drugs predicted by DFDRNN for AD. Among these, Amantadine is the first drug predicted to hold therapeutic potential for AD. Amantadine, commonly used as an antiviral and antiparkinson drug, exhibits actions as an antiviral, antiparkinson, dopaminergic, and analgesic agent. It can also be utilized for the prevention or symptomatic treatment of influenza A. Previous studies have shown improvements in the mental state of two AD patients following Amantadine treatment 44. Carbamazepine functions as an anticonvulsant and analgesic drug. Earlier studies have suggested that the combined therapy of carbamazepine and haloperidol holds promise in the clinical management of elderly AD patients 45. DFDRNN predicts an association with carbamazepine, which has been validated by CTD. Moreover, all the other drugs listed in Table 5 have been confirmed by relevant authoritative data sources, with a success rate of 100%.
Table 5: The top 10 candidate drugs for AD are ranked based on the DFDRNN prediction score